A pharmacist needs 60 oz of a 30% acid solution. He had available A which is 20% acid and B which is 50% acid. How many oz of each should be mixed to get the desired amount?
Answer by Alan3354(69443)

A pharmacist needs 60 oz of a 30% acid solution. He had available A which is 20% acid and B which is 50% acid. How many oz of each should be mixed to get the desired amount?
--------------
A + B = 60 (total of 60 oz)
0.2A + 0.5B = 0.3*60 (total of acid in solutions)
----------
B = 60-A
2A + 5B = 180
2A + 5(60-A) = 180
2A + 300 - 5A = 180
-3A = -120
A = 40 oz
B = 20 oz